# Mean estimation in the add-remove model of differential privacy

### 11 Dec 2023 v1

@@ 5 Lower bound
$R_{ar}(\varepsilon,\ell,u)=(u-\ell)^{2}R_{ar}(\varepsilon,0,1).$
@@
$L(\hat{\mu}^{\prime},D)\leq(u-\ell)^{2}L(\hat{\mu},f(D)).$
$R_{ar}(\varepsilon,\ell,u)\leq(u-\ell)^{2}R_{ar}(\varepsilon,0,1).$
The proof for the other direction is similar and omitted. ∎
We now state a well known result from Geng and Viswanath (2014), that we will use to prove the bound.
Lemma 4 ( Geng and Viswanath (2014, Section VI.C) ).
Let $k>0$ and $n>0$. Let $D_{i}$ denote the dataset that contains $i$ values of $1$ and $n$ values of 00. Let $\mathcal{D}^{k}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$. Let $N(D_{i})$ denote the number of $1$s in $D_{i}$ and let $\hat{N}$ denote an estimator. Then for any $\varepsilon<\varepsilon_{0}\ll 1$,
@@
and the best way to get $\hat{N}$ is to noise $N(D_{i})$ with a geometric mechanism. Here $o(1)$ term goes to zero as $k$ tends to infinity and $\varepsilon$ goes to 00.
Theorem 2.
$R_{ar}(\varepsilon,\ell,u)\geq\frac{2(u-l)^{2}}{\varepsilon^{2}}\left(1-o(1)\right).$
@@
By Lemma 3, it suffices to consider the scenario when $\ell=0$ and $R=1$. Let $\mathcal{D}^{k}$ and $N(D_{i})$ be the same as those be defined as in Lemma 4. Observe for any dataset $D$ in $\mathcal{D}^{k}$,
$\mu(D)=\frac{N(D)}{n+N(D)}\leq\frac{k}{n}.$
Suppose we have an $\varepsilon$-differentially private estimator $\hat{\mu}$ on $\mathcal{D}^{k}$. We convert it to an estimator of $N(D)$ as
$\hat{N}=n\hat{\mu}(D).$
$\mathbb{E}\left[(\hat{N}-N(D))^{2}\right]\geq\frac{2}{\varepsilon^{2}}(1-o(1))$
and hence
$\mathbb{E}\left[(n\hat{\mu}-N(D))^{2}\right]\geq\frac{2}{\varepsilon^{2}}(1-o(1))$
We now upper bound the left hand side of the above expression.
$\mathbb{E}\left[(n\hat{\mu}-N(D))^{2}\right]$ $=\mathbb{E}\left[(n\hat{\mu}-n\mu+n\mu-N(D))^{2}\right]$
$=\mathbb{E}\left[(n\hat{\mu}-n\mu+n\mu-N(D))^{2}\right]$
$=\mathbb{E}\left[(n\hat{\mu}-n\mu-\mu(D)N(D))^{2}\right]$
$=\mathbb{E}\left[(n\hat{\mu}-n\mu)^{2}\right]+\mathbb{E}\left[(\mu(D)N(D))^{2}\right]-2\mathbb{E}\left[(n\hat{\mu}-n\mu)(\mu(D)N(D))\right]$
where the first inequality follows by Cauchy-Schwarz inequality and the second inequality follows by observing that both $\hat{\mu}$ and $\mu$ lie in $[0,k/n]$. Setting $k=o(n^{1/3}/\varepsilon^{1/3})$ yields the following theorem. ∎
6 Experiments
(a) Normalized MSE of Algorithm 3.
(b) Ratio of Normalized MSE of Algorithm 3 to the theoretical bound for $\varepsilon=0.5$.
(c) Ratio of mean squared errors of Algorithm 3 and Algorithm 2.
Figure 2: All experiments are conducted over datasets of size $1,000$ and experiments are repeated $10,000$ times for statistical consistency.
Without loss of generality, we set $\ell=0$ and $u=1$ throughout the experiments. All datasets have $1,000$ points and the experiments are repeated $10,000$ times for statistical consistency. In Figure 1(a), we plot the performance of Algorithm 3 for different values of $\varepsilon$ and for different values of true mean. Observe that the experiments validate theory. To further verify the difficult cases for the proposed algorithm, we fix $\varepsilon=0.5$ and vary $\mu$ and plot the ratio of normalized MSE to the theoretical bound of $2/\varepsilon^{2}$ in Figure 1(b). Observe that as the theory predicts, the ratio is highest when $\mu=\Theta(1/n\varepsilon)$.
Finally, we plot the performance ratio of MSE of Algorithms 2 and 3 for various values of $\varepsilon$ and for different values of true mean in Figure 1(c). Observe that the factor is roughly around two.
7 Acknowledgements
The authors thank Travis Dick and Ziteng Sun for helpful comments and suggestions.
Appendix A Proof of Lemma 1

### 19 Feb 2024 v2

@@ Appendix A Properties of staircase mechanisms
Let $\varepsilon\geq\varepsilon_{0}$ for a sufficiently large value of $\varepsilon_{0}$. Then, for the two-dimensional stair-case mechanism,
$\min_{\gamma}\mathbb{E}_{f_{\gamma}}[x^{2}]=\min_{\gamma}\mathbb{E}_{f_{\gamma}}[y^{2}]=\Delta^{2}\cdot\Theta\left(e^{-\varepsilon/2}\right),$
@@
Without loss of generality, we assume the sensitivity to be one. By symmetry,
$\min_{\gamma}\mathbb{E}_{f_{\gamma}}[x^{2}]=\min_{\gamma}\mathbb{E}_{f_{\gamma}}[y^{2}],$
and it suffices to consider $\mathbb{E}_{f_{\gamma}}[x^{2}]$. For $k\geq 1$, it can be shown that
@@
Similarly,
[…]
and
[…]
Furthermore,
$a(\gamma)=\Theta\left(\frac{1}{e^{-\varepsilon}+\gamma^{2}}\right).$
Combining the above four equations yields,
$\mathbb{E}_{f_{\gamma}}[x^{2}]$ $=a(\gamma)\Theta\left(\gamma^{4}+e^{-\varepsilon}+\sum_{k\geq 1}(e^{-k\varepsilon}+e^{(k+1)\varepsilon})k^{3}\right)$
$=a(\gamma)\Theta\left(\gamma^{4}+e^{-\varepsilon}\right)$
$=\Theta\left(\frac{\gamma^{4}+e^{-\varepsilon}}{e^{-\varepsilon}+\gamma^{2}}\right).$
Minimizing over $\gamma$ yields the desired result. ∎
We next prove a result on the moments of both one-dimensional and two-dimensional staircase mechanisms.
Lemma 8.
Let $16\geq m\geq 0$. For both the one-dimensional and two-dimensional staircase mechanisms that guarantees $\varepsilon$-differential privacy for sensitivity one queries,
$\mathbb{E}[x^{m}]=\frac{1}{\varepsilon^{m-2}}\cdot O\left(\mathbb{E}[x^{2}]\right)$
@@
We provide the proof for the univariate staircase mechanism. The proof for the two-dimensional stair case mechanism is similar and omitted. Let $a(\gamma)$ denote the normalization term in the definition of univariate staircase mechanism Geng et al. [2015, equation 26]. Since $\gamma\leq 1$,
$\mathbb{E}_{f_{\gamma}}[x^{m}]$ $=a(\gamma)\Theta\left(\gamma^{m}+\sum_{k\geq 1}e^{-k\varepsilon}k^{m}\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}a(\gamma)O\left(\gamma^{2}+\sum_{k\geq 1}e^{-k\varepsilon}k^{m}\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}a(\gamma)O\left(\gamma^{2}+\frac{1}{(1-e^{-\varepsilon})^{m-2}}\sum_{k\geq 1}e^{-k\varepsilon}k^{2}\right)$
$=\frac{1}{(1-e^{-\varepsilon})^{m-2}}a(\gamma)O\left(\gamma^{2}+\sum_{k\geq 1}e^{-k\varepsilon}k^{2}\right)$
$=\frac{1}{(1-e^{-\varepsilon})^{m-2}}\cdot O\left(\mathbb{E}_{f_{\gamma}}[x^{2}]\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\frac{1}{\varepsilon^{m-2}}\cdot O\left(\mathbb{E}_{f_{\gamma}}[x^{2}]\right),$
where $(a)$ uses the fact that $\gamma\leq 1$, $(c)$ follows from the fact that $1-e^{-\varepsilon}\leq\varepsilon$. To prove $(b)$ notice that
$(1-e^{-\varepsilon})\sum_{k\geq 1}e^{-k\cdot\varepsilon}k^{m}=\sum_{k\geq 1}e^{-k\cdot\varepsilon}(k^{m}-(k-1)^{m})=O\left(\sum_{k\geq 1}e^{-k\cdot\varepsilon}k^{m-1}\right).$
We next state a concentration property of both one-dimensional and two-dimensional staircase mechanisms.
Lemma 9.
Let $n\geq 10$. For both the one-dimensional and two-dimensional staircase mechanisms that guarantees $\varepsilon$-differential privacy for sensitivity one,
[…]
Proof.
We provide the proof for the univariate staircase mechanism. The proof for the two-dimensional stair case mechanism is similar and omitted. Let $a(\gamma)$ denote the normalization term in the definition of univariate staircase mechanism Geng et al. [2015, equation 26]. Note that for the univariate staircase mechanism,
$a(\gamma)\leq\frac{1-e^{-\varepsilon}}{2e^{-\varepsilon}}.$
We now bound the desired quantity as follows.
$\text{Pr}(x\leq-n)$ $\leq\sum_{k\geq n}a(\gamma)e^{-k\varepsilon}$
$=a(\gamma)\frac{e^{-n\varepsilon}}{1-e^{-\varepsilon}}$
$\leq(1-e^{-\varepsilon})e^{\varepsilon}\frac{e^{-n\varepsilon}}{1-e^{-\varepsilon}}$
$=e^{-(n-1)\varepsilon},$
where the last inequality follows from the bound on $a(\gamma)$. ∎
Appendix B Proof of Lemma 1
@@
Appendix C Proof of Lemma 2
$\mathbb{E}[C^{2}]$ $=\mathbb{E}\left[\left(\frac{Z_{s}}{n}-\frac{(\mu-m)Z_{n}}{n}\right)^{2}\right]$
$=\frac{2(u-l)^{2}}{n^{2}\varepsilon^{2}}+\frac{8(\mu-m)^{2}}{n^{2}\varepsilon^{2}}.$
Since $(x+y)^{2}\leq 2x^{2}+2y^{2}$,
$\mathbb{E}[F^{2}]$ $\leq 8\frac{(u-\ell)^{2}}{n^{4}}\mathbb{E}\left[Z^{4}_{n}\right]+8\frac{1}{n^{4}}\mathbb{E}\left[Z^{2}_{a}Z^{2}_{b}\right]$
$=o\left(\frac{(u-l)^{2}}{n^{2}\varepsilon^{2}}\right).$
Finally, by the tail bounds of the Laplace mechanism,
$M^{2}\text{Pr}(Z_{b}<-b/2)$ $\leq(u-\ell)^{2}\text{Pr}(Z_{n}<-n/2)$
$=o\left(\frac{(u-l)^{2}}{n^{2}\varepsilon^{2}}\right).$
Combining the above three equations together with Lemma 1 yields the lemma.
Appendix D Analysis of the Hourglass Mechanism
D.1 Computing normalization constant
Let $c(\gamma)$ be the normalizing factor that ensures the sum of the probabilities is one.
From the paper Geng and Viswanath [2014], we know the normalizing factor for the univariate staircase mechanism to be
$a(\gamma)\triangleq\frac{1-e^{-\varepsilon}}{2(\gamma+^{-\varepsilon}(1-\gamma))}.$
From Lemma 3, the marginal of the hourglass distribution is the univariate staircase distribution. hence,
$\sum_{y}f_{\gamma}(x,y)=a(\gamma),$
where $a(\gamma)$ is the normalization factor in the univariate staircase distribution. Observe that if $x=0$, then for all integer $y$, $(x,y)\in A_{y}$ and hence
$\sum_{y}f_{\gamma}(x,y)=\sum_{y}c(\gamma)e^{-\varepsilon y}=c(\gamma)\frac{1-e^{-\varepsilon}}{1+e^{-\varepsilon}}.$
Hence,
$c(\gamma)=\frac{1-e^{-\varepsilon}}{1+e^{-\varepsilon}}a(\gamma)=\frac{(1-e^{-\varepsilon})^{2}}{2\Delta(1+e^{-\varepsilon})(\gamma+e^{-\varepsilon}(1-\gamma))}.$
D.2 Proof of Theorem 2
Let $D$ and $D^{\prime}$ be two neighboring datasets such that $q(D)-Q(D_{0})=-(x_{0},1-x_{0})$. Let let $\tilde{q}(D)$ denote the output of the hourglass mechanism. To provide differential privacy guarantee, it suffices to prove upper and lower bounds for
$\frac{\text{Pr}(\tilde{q}(D)=(x,y))}{\text{Pr}(\tilde{q}(D^{\prime})=(x,y))}.$
Let $\bar{Z}$ be a sample from the hourglass mechanism, then
$\frac{\text{Pr}(\tilde{q}(D)=(x,y))}{\text{Pr}(\tilde{q}(D^{\prime})=(x,y))}$ $=\frac{\text{Pr}({q}(D)+\bar{Z}=(x,y))}{\text{Pr}({q}(D^{\prime})+\bar{Z}=(x,y))}$
$=\frac{\text{Pr}({q}(D^{\prime})-(x_{0},1-x_{0})+\bar{Z}=(x,y))}{\text{Pr}({q}(D^{\prime})+\bar{Z}=(x,y))}$
$=\frac{\text{Pr}(\bar{Z}=(x,y)-{q}(D^{\prime})+(x_{0},1-x_{0}))}{\text{Pr}(\bar{Z}=(x,y)-{q}(D^{\prime}))}$
$=\frac{\text{Pr}(\bar{Z}=(x^{\prime},y^{\prime})+(x_{0},1-x_{0}))}{\text{Pr}(\bar{Z}=(x^{\prime},y^{\prime})}$
$=\frac{f_{\gamma}(x^{\prime}+x_{0},y^{\prime}+1-x_{0}))}{f_{\gamma}(x^{\prime},y^{\prime})},$
where $(x^{\prime},y^{\prime})=(x,y)-{q}(D^{\prime})$. Hence, to prove the privacy guarantee, it suffices to prove upper and lower bounds on the ratio,
$R(x,y,x_{0})\triangleq\frac{f_{\gamma}(x+x_{0},y+1-x_{0}))}{f_{\gamma}(x,y)},$
for all $x,y$ and $x_{0}\in[0,1]$. Without loss of generality, we assume that $\Delta=1$. Let $\mathcal{S}=\{A_{k},B_{k}(i):\forall k,i\}$. Note that $\mathcal{S}$ partitions the domain of $f_{\gamma}$ into disjoint partitions. We observe that if $(x,y)\in A_{-1}$, then $f_{\gamma}(x,y)=c(\gamma)e^{-\varepsilon}$ and for all $x_{0}$, $f_{\gamma}(x+x_{0},y+1-x_{0})\in\{c(\gamma),c(\gamma)e^{-2\varepsilon}\}$, hence if $(x,y)\in A_{-1}$ then $R(x,y,x_{0})\in\{e^{-\varepsilon},e^{\varepsilon}\}$ for all $x_{0}$. If $(x,y)\notin A_{-1}$, then $(x+x_{0},y+1-x_{0})$ belongs to at most two sets in $\mathcal{S}$ and hence $f_{\gamma}(x+x_{0},y+1-x_{0})$ is monotonic in $x_{0}$ and proving the result for $x_{0}\in\{0,1\}$ suffices.
We now focus on the case when $x_{0}\in\{0,1\}$. Of these two cases, by symmetry it suffices to consider $x_{0}=0$. Furthermore, we show the result when $x+y=k$ for some $x\geq 0$. The proof for the other side is similar and omitted. We prove the result by dividing the problem into subcases depending the value of $(x,y)$. Subcase $(A)$: If $(x,y)\in A_{k}$ for some $k$, then $(x,y+1)\in A_{k+1}$ and hence $R(x,y,0)=e^{-\varepsilon}$. Subcase $(B1)$: If $(x,y)\in B_{k}(1)$, then $(x,y)\in A_{k+1}$ and hence $R(x,y,0)=e^{\varepsilon}$. Subcase $(B2)$: If $(x,y)\in B_{k}(i)$ for $i\geq 2$, then $(x,y)\in B_{k+1}(i-1)$ and hence $R(x,y,0)=e^{\varepsilon}$. Hence, we have shown that for all $x,y$ and $x_{0}\in\{0,1\}$,
$R(x,y,x_{0})\in\{e^{-\varepsilon},e^{\varepsilon}\}.$
Appendix E Proof of Lemma 5
The privacy guarantee is similar to that of Theorem 1 and is omitted. As before, the proof of utility heavily relies in Lemma 1. Let $n=|D|$. Observe that
$\mathbb{E}[(\hat{\mu}-\mu)^{2}]\leq(u-\ell)^{2}\mathbb{E}\left[\left(\frac{\hat{s}_{1}}{\hat{s}_{1}+\hat{s}_{2}}-\frac{s_{1}}{n}\right)^{2}\right].$
Let $a=s_{1}$, $b=n$, $Z_{a}=Z_{1}$, $Z_{b}=Z_{1}+Z_{2}$, where $Z_{1},Z_{2}$ are from the two-dimensional staircase mechanism. Let $M=1$. With these definitions, to apply Lemma 1, we need to bound $\mathbb{E}[C^{2}]$, $\mathbb{E}[F^{2}]$, and $\text{Pr}(Z_{b}<-b/2)$. We bound each of the terms below. Let $\alpha=\frac{s_{1}}{n}$.
$\mathbb{E}[C^{2}]$
$=\mathbb{E}\left[\left(\frac{Z_{1}}{n}-\frac{(\frac{s_{1}}{n})(Z_{1}+Z_{2})}{n}\right)^{2}\right]$
$=\frac{1}{n^{2}}\mathbb{E}\left[(1-\alpha)^{2}Z^{2}_{1}+\alpha^{2}Z^{2}_{2}-2(1-\alpha)\alpha Z_{1},Z_{2}\right]$
$\leq\frac{1}{n^{2}}\mathbb{E}\left[(1-\alpha)^{2}Z^{2}_{1}+\alpha^{2}Z^{2}_{2}\right]+2(1-\alpha)\alpha\sqrt{\mathbb{E}\left[Z^{2}_{1},Z^{2}_{2}\right]}$
$\leq\frac{1}{n^{2}}(1-\alpha)^{2}\widetilde{\sigma}^{2}(\varepsilon)+\alpha^{2}\widetilde{\sigma}^{2}(\varepsilon)+2(1-\alpha)\alpha\widetilde{\sigma}^{2}(\varepsilon)$
$=\frac{\widetilde{\sigma}^{2}(\varepsilon)}{n^{2}},$
where the first inequality follows by the Cauchy-Schwarz inequality. Since $(x+y)^{2}\leq 2x^{2}+2y^{2}$,
$\mathbb{E}[F^{2}]$ $\leq\frac{8}{n^{4}}\mathbb{E}[(Z_{1}+Z_{2})^{4}]+\frac{8}{n^{4}}\mathbb{E}[(Z_{1}+Z_{2})^{2}Z^{2}_{1}]$
$=o\left(\frac{\widetilde{\sigma}^{2}(\varepsilon)}{n^{2}}\right),$
where the last equality follows from Lemma 8. Finally,
$\text{Pr}(Z_{b}<-b/2)$ $\leq\text{Pr}(Z_{1}+Z_{2}<-n/2)$
$\leq\text{Pr}(Z_{1}<-n/4)+\text{Pr}(Z_{2}<-n/4)$
$=o\left(\frac{\widetilde{\sigma}^{2}(\varepsilon)}{n^{2}}\right),$
where the last equality follows from Lemma 9. Combining the above three equations together with Lemma 1 and observing the fact that $\mu=\ell+\frac{s_{1}w}{n}$ yields the result.
Appendix F Proof of Theorem 3
The privacy guarantee is similar to that of Theorem 1 and is omitted. As before, the proof of utility heavily relies in Lemma 1. Let $n=|D|$. Observe that
$\mathbb{E}[(\hat{\mu}-\mu)^{2}]\leq(u-\ell)^{2}\mathbb{E}\left[\left(\frac{\hat{s}_{1}}{\hat{s}_{1}+\hat{s}_{2}}-\frac{s_{1}}{n}\right)^{2}\right].$
Let $a=s_{1}$, $b=n$, $Z_{a}=Z_{1}$, $Z_{b}=Z_{1}+Z_{2}$, where $Z_{1},Z_{2}$ are from the hourglass mechanism. Let $M=1$. With these definitions, to apply Lemma 1, we need to bound $\mathbb{E}[C^{2}]$, $\mathbb{E}[F^{2}]$, and $\text{Pr}(Z_{b}<-b/2)$. We bound each of the terms below. Let $\alpha=\frac{s_{1}}{n}$.
$\mathbb{E}[C^{2}]$
$=\mathbb{E}\left[\left(\frac{Z_{1}}{n}-\frac{(\frac{s_{1}}{n})(Z_{1}+Z_{2})}{n}\right)^{2}\right]$
$=\frac{1}{n^{2}}\mathbb{E}\left[(1-\alpha)^{2}Z^{2}_{1}+\alpha^{2}Z^{2}_{2}-2(1-\alpha)\alpha Z_{1},Z_{2}\right]$
$\leq\frac{1}{n^{2}}\mathbb{E}\left[(1-\alpha)^{2}Z^{2}_{1}+\alpha^{2}Z^{2}_{2}\right]+2(1-\alpha)\alpha\sqrt{\mathbb{E}\left[Z^{2}_{1},Z^{2}_{2}\right]}$
$\leq\frac{1}{n^{2}}(1-\alpha)^{2}\widetilde{\sigma}^{2}(\varepsilon)+\alpha^{2}\widetilde{\sigma}^{2}(\varepsilon)+2(1-\alpha)\alpha\widetilde{\sigma}^{2}(\varepsilon)$
$=\frac{{\sigma}^{2}(\varepsilon)}{n^{2}},$
where the first inequality follows by the Cauchy-Schwarz inequality. Since $(x+y)^{2}\leq 2x^{2}+2y^{2}$,
$\mathbb{E}[F^{2}]$ $\leq\frac{8}{n^{4}}\mathbb{E}[(Z_{1}+Z_{2})^{4}]+\frac{8}{n^{4}}\mathbb{E}[(Z_{1}+Z_{2})^{2}Z^{2}_{1}]$
$=o\left(\frac{{\sigma}^{2}(\varepsilon)}{n^{2}}\right),$
where the last equality follows from Lemma 8 and the fact that the marginal distribution of hourglass mechanism is same as the staircase mechanism.. Finally,
$\text{Pr}(Z_{b}<-b/2)$ $\leq\text{Pr}(Z_{1}+Z_{2}<-n/2)$
$\leq\text{Pr}(Z_{1}<-n/4)+\text{Pr}(Z_{2}<-n/4)$
$=o\left(\frac{{\sigma}^{2}(\varepsilon)}{n^{2}}\right),$
where the last equality follows from Lemma 9 and the fact that the marginal distribution of hourglass mechanism is same as the staircase mechanism. Combining the above three equations together with Lemma 1 and observing the fact that $\mu=\ell+\frac{s_{1}w}{n}$ yields the theorem.
Appendix G Proof of Theorem 4
Lemma 10.
[…]
Proof.
$L(\hat{\mu}^{\prime},D)\leq(u-\ell)^{2}L(\hat{\mu},f(D)).$
$R_{\text{ar}}(\varepsilon,\ell,u)\leq(u-\ell)^{2}R_{\text{ar}}(\varepsilon,0,1).$
The proof for the other direction is similar and omitted. ∎
Proof of Theorem 4.
By Lemma 10, it suffices to consider the scenario when $\ell=0$ and $u=1$. Let $\mathcal{D}^{k}$ and $S(D)$ be the same as those be defined as in Lemma 6. Let $\widetilde{\mathcal{D}}^{k}$ be the set of all datasets obtaining by combining each dataset in $\mathcal{D}^{k}$ with $n$ values of ones. Observe for any dataset $D$ in $\widetilde{\mathcal{D}}^{k}$,
$\mu(D)=\frac{S(D)}{n+k}\leq\frac{k}{n}.$
Suppose we have an $\varepsilon$-differentially private estimator $\hat{\mu}$ on $\mathcal{D}^{k}$. We convert it to an estimator of $S(D)$ as
$\hat{S}=n\hat{\mu}(D).$
$\mathbb{E}\left[(\hat{S}-S(D))^{2}\right]\geq\frac{2}{\varepsilon^{2}}(1-o(1))$
and hence
$\mathbb{E}\left[n\hat{\mu}-S(D))^{2}\right]\geq\frac{2}{\varepsilon^{2}}(1-o(1)),$
which implies
$|D|^{2}\mathbb{E}\left[n^{2}\hat{\mu}-S(D))^{2}\right]\geq\frac{2}{\varepsilon^{2}}(1-o(1)).$
We now upper bound the left hand side of the above expression.
$\mathbb{E}\left[(n\hat{\mu}-S(D))^{2}\right]$
$=\mathbb{E}\left[(n\hat{\mu}-n\mu+n\mu-S(D))^{2}\right]$
$=\mathbb{E}\left[(n\hat{\mu}-n\mu+n\mu-S(D))^{2}\right]$
$=\mathbb{E}\left[(n\hat{\mu}-n\mu-\mu(D)k)^{2}\right]$
$=\mathbb{E}\left[(n\hat{\mu}-n\mu)^{2}\right]+\mathbb{E}\left[(\mu(D)k)^{2}\right]$
$-2\mathbb{E}\left[(n\hat{\mu}-n\mu)(\mu(D)k)\right]$
where the first inequality follows by Cauchy-Schwarz inequality and the second inequality follows by observing that both $\hat{\mu}$ and $\mu$ lie in $[0,k/n]$. Setting $k=o(n^{1/3})$ yields the following theorem. ∎
Appendix H Additional experiments
Figure 3: hourglass and two dimensional staircase mechanisms, low privacy regime, varying $\mu$.
In this section, we explore how the MSEs of different algorithms change when the the size of the dataset $|D|$ changes. The errors are normalized by $\frac{\varepsilon^{2}}{2}$ to keep them in a similar range across different privacy regimes. We choose $|D|\in\{2^{i}\}_{i=0,\ldots,6}$ and observe the trend of changes in MSEs for all listed algorithms.
[…]
Figure H shows how the MSEs (normalized by $\frac{\varepsilon^{2}}{2}$) of different Laplace mechanisms vary as a function of $|D|$ for $\varepsilon=1.0$ and $\mu=0.01$. The lower bound (red curve), as in Section 6, is developed from using the univariate staircase mechanism with optimal $\gamma$ on private mean in the swap model. One can see that only Algorithm 3 converges to the lower bound as $|D|$ grows bigger, being about two times better than Algorithm 2 when $|D|\geq 16$.
On the other hand, Figure H and H compare the MSEs (normalized by $\frac{\varepsilon^{2}}{2}$) of the two-dimensional staircase and the hourglass mechanism in the low privacy regime when $\varepsilon=4.0$ and $\mu=0.01$. The ratio of the former over latter is found to be always bigger than one, showing hourglass mechanism to be better.